OpenStudy (anonymous):

if a ball hit a wall at velocity of v then by knowing the P(1)=P(2) why there is no kinetic energy at the wall?

OpenStudy (anonymous):

Some of the kinetic energy would have been transferred into other forms of energy - the sound of the impact, the deformation of the ball and its rebounding away from the wall, possibly some heat, etc.

OpenStudy (anonymous):

During collisions, momentum is always conserved, while energy may not be. Momentum is expressed as\[p = mv\]and kinetic energy is expressed as\[KE = {1 \over 2} m v^2\]Since the momentum of the ball before and after collision are equal (we can assume the mass of the ball is constant) we can say that \[v_1 = v_2\]This being said, the wall has zero velocity. From the expression for kinetic energy, if velocity is zero the kinetic energy must also be zero.

OpenStudy (anonymous):

v do know that P(1)=P(2) so M(ball)V(ball)=M(wall)*V(wall) so the wall must have a velocity which can be calculate from the following equation and we do know too that\[K=1/2M(wall)*V^2(wall)\] we have the M and we have V but there is no kinetic energy WHY?

OpenStudy (anonymous):

Not quite. eashmore explained it quite well (ignore my prior explanation, I misunderstood your question). Does the wall generally move if you throw a small ball at it?

OpenStudy (anonymous):

yea/because there is a force from the ball \[F=m*a\] so there is an acceleration and we can write down : \[V^2-V^2=2a \Delta x\] so the wall moves.

OpenStudy (anonymous):

No, it doesn't. Or if it does, it's immeasurably small and not worth consideration. The mass of the ball is very small in comparison to the mass of the wall, so any kinetic energy the ball has is going to be greatly insufficient to overcome the wall's inertia. It's like if I, on foot, ran as fast I could headfirst into a stationary jumbo jet - the jet wouldn't be the one bouncing off me.

OpenStudy (anonymous):

just because the wall provides a force back on the ball doesn't mean there is motion. The force onto the ball from the wall is always equal and opposite to whatever the ball generates. This is assuming we aren't talking about a wrecking ball or something like that. It's the same as you sitting in your chair. You feel a push back up from the chir. If all of a sudden someone gave you a 50lb weight, you would fell more push up from your chair. Equal and opposite. This is assuming the chair is structurally capable of supporting the weight.

OpenStudy (anonymous):

can anyone answer , use the ideal gas law to explain the changes in temperature, pressure and volume of the coolant gas at the compressors stage of its refrigeration

OpenStudy (anonymous):

Please don't post your question in other threads esmail, someone will get to answering it eventually.

OpenStudy (anonymous):

ok but can you guys please try and answer this

OpenStudy (anonymous):

\[P(1)V(1)/T(1)=P(2)V(2)/T(2)\] use the followings equations to find out ur answer...